OpenStudy (anonymous):

find the total cost of making a triangular sail that has a base dimension of 14 feet and a height of 30 feet if the price for making the sail is $20.50 per square foot. what is the total cost of making the triangular sail?

OpenStudy (aaliyahfairgood):

first find the area of the triangle A=1/2 (base times height) Area= 1/2 (14 times 30) Area= 1/2 (420) Area=210 ft squared Then multiply 210 ft by the cost per square foot 210 times $20.50 =$4305